The combined cost of one advance ticket to a show and one same-day ticket was $55. It is known that 30 tickets were sold in advance and 25 the same day, for total receipts of $1575 . What was the price of each kind of ticket?

Answer by Paul(988)

Let the cost of the Advance ticket be x
Let the cost of the same day ticket be y
x+y=55
y=55-x (subsituttion)
30x+25y=1575
30x+25(55-x)=1575
5x=200
x=40
y=55-40
y=15
Hence, the advance ticket costs $40 and the same day ticket costs $15.
Paul.